OpenStudy (bananas1234):

Simplify: square root x^14

geerky42 (geerky42):

You just use this rule \((a^b)^c = a^{bc}\)

geerky42 (geerky42):

\[\sqrt{x^{14}} = (x^{14})^{\frac12}\]

geerky42 (geerky42):

\[(x^{14})^{\frac12} = x^{14\cdot\frac12}\]

OpenStudy (bananas1234):

x^7?

OpenStudy (bananas1234):

@geerky42

geerky42 (geerky42):

Yep.

OpenStudy (bananas1234):

Thank you, i had posted this question before but i was still confused and no one was answering so i re-posted, idk if you are supposed to do that. Hope that was ok :) @geerky42

geerky42 (geerky42):

It's cool.